# Supervised Learning Models for Early Detection of Albuminuria Risk in Type-2 Diabetes Mellitus Patients

### v1

@@ I Introduction
II Methodology
II-A Dataset
@@
II-B Design of Experiment
This study aimed to evaluate the performance of supervised learning algorithms in predicting the risk of developing albuminuria in patients with T2DM patients. We evaluated several supervised learning algorithms, including 6 machine learning algorithms and 1 deep learning algorithm. The machine learning algorithms used were Naïve Bayes, Support Vector Machine (SVM), decision tree, random forest, AdaBoost, and XGBoost. Among these machine learning algorithms, random forest, AdaBoost, and XGBoost are ensemble algorithms. The deep learning algorithm employed in this study is the Multi-Layer Perceptron (MLP). Table II presents the complete experimental design used in this study. We used the scikit-learn library [12], version 1.0.2, as our primary machine learning and deep learning toolkit. Additionally, we employed the xgboost library [13], version 1.6.2, specifically designed for implementing the XGBoost algorithm.
@@ III Result
The superior results were obtained from the deep learning algorithm, specifically the Multi-Layer Perceptron (MLP), which achieved an accuracy and f1-score of 0.74 and 0.71, respectively. This algorithm outperformed the machine learning algorithms, which only achieved accuracy and f1-scores of up to 0.65 and 0.55, respectively. Additionally, the MLP algorithm exhibited the highest precision and recall scores compared to the other algorithms, scoring 0.68 and 0.75, respectively. This result indicates that the algorithm might be acceptable for predicting the risk of albuminuria among T2DM patients. However, further improvements are needed, particularly in terms of the dataset size and variety, to achieve better results.
@@
Despite the complex nature of the patient data, the outperforming performance of the MLP algorithm might be beneficial due to its architecture. The MLP consists of several to even thousands of hidden layers. The MLP uses the backpropagation algorithm to update its weights based on the learned data [14]. This enables the MLP to solve complex problems relatively easily compared to other traditional machine learning algorithms when dealing with such complex data.
IV Conclusion

### v4

@@ I Introduction
II Related Work
Recently, there has been a growing interest among researchers in using machine learning approaches to predict albuminuria. This interest arise from the urgency of developing early risk prediction tools for the disease, as it can lead to increased ”costs” if left undetected. To our knowledge, two studies conducted by Khitan et al. [8] and Lin et al. [9] have used machine learning approaches for predicting albuminuria.
Khitan et al. [8] in their study used machine learning approaches to predict the risk of albuminuria in person with diabetes. Their study incorporated 13 predictive factors, including measures such as subtotal lean mass, subtotal fat mass, diabetes duration, age, HbA1c levels, creatinine levels, triglyceride levels, total cholesterol levels, HDL cholesterol levels, maximum exercise capacity, systolic and diastolic blood pressure, and ankle brachial index. They conducted their study on 1330 subjects and used a variety of machine learning algorithms, including random forest, gradient boost, logistic regression, support vector machines, multilayer perceptron, and a stacking classifier. The results showed that the multilayer perceptron (MLP) exhibited the highest performance with an AUC (Area Under the Curve) value of 0.67. Furthermore, the model demonstrated a precision of 0.61, recall of 0.67, and an accuracy of 0.62, as determined from the confusion matrix presented in the paper.
In another study within this domain, Lin et al. [9] aimed to predict microalbuminuria in the Chinese population using machine learning approaches. Their study involved 3,294 subjects ranging in age from 16 to 93 years. They used the ”glm” package in the R software to construct their machine learning model. Their model achieved a specificity of 0.9 and an accuracy of 0.63, although the sensitivity was relatively low at 0.2. Despite these outcomes, the study’s conclusions highlighted systolic and diastolic blood pressure, fasting blood glucose levels, triglyceride levels, gender, age, and smoking as potential predictors of microalbuminuria among the patient population.
III Methodology
III-A Dataset
@@
Figure 2 illustrates the simplified mechanism of diabetic nephropathy, with obesity playing a central role. Elevated BMI increases the likelihood of obesity, which subsequently acts as a risk factor for developing diabetes and hypertension through a complex pathway. The intricate sequence involves the increase of plasma glucose and HbA1c in diabetes, leading to microvascular damage and subsequent kidney damage. The duration of diabetes increases the risk of such damage. On the other hand, chronic high blood pressure resulting from obesity can lead to hypertension, causing microvascular damage and putting the individual at risk of kidney damage. Additionally, kidney damage can, in turn, induce hypertension, creating an inner loop-like mechanism that worsens the condition. Furthermore, obesity also serves as a risk factor for lipid profile issues, such as an increase in LDL, TG, and cholesterol, and a decrease in HDL, posing a risk of dyslipidemia. Dyslipidemia, in turn, indirectly contributes to the kidney damage.
@@
III-B Design of Experiment
This study aimed to evaluate the performance of supervised learning algorithms in predicting the risk of developing albuminuria in patients with T2DM patients. We evaluated several supervised learning algorithms, including 6 machine learning algorithms and 1 deep learning algorithm. The machine learning algorithms used were Naïve Bayes, Support Vector Machine (SVM), decision tree, random forest, AdaBoost, and XGBoost. Among these machine learning algorithms, random forest, AdaBoost, and XGBoost are ensemble algorithms. The deep learning algorithm employed in this study is the Multi-Layer Perceptron (MLP). The use of deep learning in this experimental design, as compared to machine learning algorithms, was intended to evaluate its potential performance considering the limited size of the dataset. Table II presents the complete experimental design used in this study. We used the scikit-learn library [14], version 1.0.2, as our primary machine learning and deep learning toolkit. Additionally, we employed the xgboost library [15], version 1.6.2, specifically designed for implementing the XGBoost algorithm.
@@ IV Result
The superior results were obtained from the deep learning algorithm, specifically the Multi-Layer Perceptron (MLP), which achieved an accuracy and f1-score of 0.74 and 0.71, respectively. This algorithm outperformed the machine learning algorithms, which only achieved accuracy and f1-scores of up to 0.65 and 0.55, respectively. Additionally, the MLP algorithm exhibited the highest precision and recall scores compared to the other algorithms, scoring 0.68 and 0.75, respectively. This result outperformed the study by Khitan et al. [8] and Lin et al. [9], indicating that the algorithm might be acceptable for predicting the risk of albuminuria among T2DM patients. However, further improvements are needed, particularly in terms of the dataset size and variety, to achieve better results.
@@
Despite the complex nature of the patient data, the outperforming performance of the MLP algorithm might be beneficial due to its architecture. The MLP might consist of several to even thousands of hidden layers. The MLP uses the backpropagation algorithm to update its weights based on the learned data [16]. This enables the MLP to solve complex problems relatively easily compared to other traditional machine learning algorithms when dealing with such complex data.
V Conclusion
@@
Acknowledgment
This study was conducted as part of the ID-CALDERA (Indonesia’s Cardiovascular Risk Stratification in Diabetes Using Smartphone-Based Retinal Imaging) project, which aims to develop AI-based models and technologies for predicting diabetic complications at an earlier stage. The vision and mission of ID-CALDERA are centered on reducing the impact and burden of diabetic complications. We extend our gratitude to everyone who has provided support for our work, whether individually or institutionally.